OpenStudy (anonymous):

PLEAASEE HELP...will medal and fan...Let f(x) = x2 + x + 2 and g(x) = 2x2 + 5. Find f(g(x)).

OpenStudy (chrisdbest):

Is that \[f(x) = x^2 + x + 2; g(x) = 2x^2 + 5\]

OpenStudy (anonymous):

yess

OpenStudy (chrisdbest):

So you have to put g(x) in for f(x)

OpenStudy (anonymous):

how

OpenStudy (chrisdbest):

So that will be \[(2x^2 + 5)^2 + (2x^2 + 5) + 2 \]

OpenStudy (chrisdbest):

Because you will be inputting every g(x) value into every x in the f(x) field.

OpenStudy (anonymous):

so now do I combine like terms or?

OpenStudy (chrisdbest):

So \[(4x^4 + 25) + (2x^2 + 5) + 2\]

OpenStudy (chrisdbest):

Yes, you can combine like terms now

OpenStudy (anonymous):

why did the 2 stay?

OpenStudy (chrisdbest):

The 2 stayed because we were plugging in g(x) terms in for f(x)

OpenStudy (chrisdbest):

So your final answer should look like \[4x^4 + 2x^2 +32\]

OpenStudy (chrisdbest):

Because the "2" didn't have a variable next to it, we didn't have to mess with it.

OpenStudy (anonymous):

ohh I see... thank you so much

OpenStudy (chrisdbest):

Does that Help?

OpenStudy (anonymous):

yes..but im very bad at it though

OpenStudy (chrisdbest):

I feel yah man, I need help in Pre-Calc because there are some parts in Pre-Calc that I go str8 to OS for. Even though I'm taking Pre-Calc online, I'm getting 90% of my answers from OS.

OpenStudy (anonymous):

I feel you..wish I could help you too. I am taking algebra online but try this website.. mathway.com

OpenStudy (anonymous):

@chrisdbest \[(2x^2+5)^2 \neq 4x^4+25\] you have to foil that \[(2x^2+5)^2 = (2x^2+5)(2x^2+5)\]

OpenStudy (chrisdbest):

Lols

OpenStudy (chrisdbest):

So \[4x^4 + 10x^2 + 25 ?\] That's what you meant?

OpenStudy (anonymous):

so confused

OpenStudy (chrisdbest):

Did you get the right answer @ sharlin?

OpenStudy (anonymous):

+20x² in the middle

OpenStudy (chrisdbest):

@sharlin ?

OpenStudy (chrisdbest):

Lols thats what I meant to pput. I'm dyslecix

OpenStudy (chrisdbest):

* put; * dyslexic

OpenStudy (anonymous):

im not sure because its not multiple choice

OpenStudy (freckles):

@peachpi was just talking about how (a+b)^2=a^2+2ab+b^2 instead of a^2+b^2 so changes will need to be made to your previous answer @sharlin

OpenStudy (anonymous):

what would the final answer be?

OpenStudy (chrisdbest):

\[4x^4 +20x^2 +25\]

OpenStudy (anonymous):

Im stil confused on how you got there

OpenStudy (anonymous):

He made the substitution and got \[f(g(x))=(2x^2+5)^2+(2x^2+5)+2\] which is the same as \[f(g(x))=(2x^2+5)(2x^2+5)+(2x^2+5)+2\] He multiplied and then combined like terms

OpenStudy (chrisdbest):

Peachpi got this. I feel like he can explain it to you better than I can.

OpenStudy (anonymous):

with me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. now multiply (2x² +5)(2x² + 5) either using foil or the rule @freckles put up

OpenStudy (anonymous):

so I got 4x^2+10x^2+10x^2+25

OpenStudy (anonymous):

Should be 4x^4. so that was just the first part of the equation. You need to combine it with the rest and combine like terms 4x^4+10x^2+10x^2+25 + 2x^2+ 5 +2

OpenStudy (anonymous):

I get a crazy answer

OpenStudy (chrisdbest):

Yo @peachpi Ima give you dat gold medal cuz I think you deserve it, cuz I don't got a lot of time to answer everyone's questions.

OpenStudy (anonymous):

@sharlin what did you get?

OpenStudy (anonymous):

thx @chrisdbest

OpenStudy (anonymous):

4x4+22x2+32..idk

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

really..im really not sure

OpenStudy (anonymous):

yep. trust your work